The following is a December 31, 2018, post-closing trial balance for Culver City Lighting, Inc. Account Title Debits Credits Cash $ 64,000 Accounts receivable 48,000 Inventories 54,000 Prepaid insurance 24,000 Equipment 120,000 Accumulated depreciation—equipment $ 43,000 Patent, net 49,000 Accounts payable 16,500 Interest payable 6,500 Note payable (due in 10, equal annual installments) 150,000 Common stock 79,000 Retained earnings 64,000 Totals $ 359,000 $ 359,000 a. Calculate the current ratio. b. Calculate the acid-test ratio. c. Calculate the debt to equity ratio.
Answer #1

a. Calculation of current ration

Current ratio = Current Assets/Current Liabilities

Current assets = 64,000+48,000+54,000+24,000 = 19,000

Current liabilities = 16,500+6,500+15,000 = 38,000

Current ratio = 190,000/38,000 = 5

b. Calculation of Acid test ratio

Acid test ratio = Quick assets/Quick liabilities

Quick assets = Total current assets - (Inventories+preliminary expenses)

= 190,000-54,000-24,000 = 112,000

Acid test ratio = 112,000/38,000 2.95

c. Calculation of

Debt to equity ratio = Debt/Equity

= (16,500+6,500+150,000)/ (79,000+64,000)

= 173,000/143,000

=1.2098(rounded)